# Spectral mechanism and nearly reducible transfer matrices for pseudotransitions in one-dimensional systems

### v1

@@ 2.2 Block-diagonal reducible matrices
This simple two-block decomposition captures the essential mechanism by which block-diagonal structure allows genuine thermodynamic singularities. In more complex cases with many disconnected sectors, the leading behavior is still often governed by the two most dominant blocks. Such effective two-block approximations are successfully applied in models like the Kittel model [43] and will be invoked again in the analysis of the nanowire model in Sec.  5.
@@ 2.3 Nearly block-diagonal irreducible matrices
$\mathbf{V}=\begin{pmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{C}^{{\rm T}}&\mathbf{B}\end{pmatrix}$, (7)
where $\mathbf{A}$ and $\mathbf{B}$ are dominant irreducible blocks and $\mathbf{C}$ contains small off-diagonal couplings. Although $\mathbf{V}$ is formally irreducible, the elements of $\mathbf{C}$ may become suppressed at a given parameter value $x_{p}$, nearly decoupling the sectors $\mathbf{A}$ and $\mathbf{B}$ at a fixed temperature
@@
$\mathcal{V}_{t}=\begin{pmatrix}w_{a}&\langle\phi_{a}|\mathbf{C}|\psi_{b}\rangle\\ \langle\phi_{b}|\mathbf{C}^{{\rm T}}|\psi_{a}\rangle&w_{b}\end{pmatrix}$, (8)
where $\langle\phi_{a}|\mathbf{C}|\psi_{b}\rangle>0$ and $\langle\phi_{b}|\mathbf{C}^{{\rm T}}|\psi_{a}\rangle>0$, according to the condition of matrix non-negative and irreducibility, then let us define $\langle\phi_{b}|\mathbf{C}^{{\rm T}}|\psi_{a}\rangle\langle\phi_{a}|\mathbf{C}|\psi_{b}\rangle=w_{0}^{2}$. Thus, the eigenvalues of $\mathcal{V}_{t}$ are:
@@ 3.1 Expansion of the spectral gap and pseudo-critical point
To analyze the behavior of thermodynamic observables near a pseudotransition, we expand the normalized spectral gap $\zeta(x,T)$ in a Taylor series around the anomalous point $(x_{p},T_{p})$, as justified by Kato’s theorem [42]. The full derivative expressions, including mixed and second-order terms, are derived in Appendix 7. Importantly, this expansion remains well-defined even when $w_{a}=w_{b}$, since the coupling $w_{0}>0$ ensures $\zeta$ stays analytic and nonzero. Expanding around the anomalous point $(x_{p},T_{p})$, we write:
@@
where we are assuming $\delta T=T-T_{p}$ and $\delta x=x-x_{p}$. Derivatives such as $\zeta_{T}=\frac{\partial\zeta}{\partial T}$, and $\zeta_{T^{2}}=\frac{\partial^{2}\zeta}{\partial T^{2}}$, are defined in the usual way. Similar definitions apply for derivatives with respect to $x$, and the mixed derivative is given by $\zeta_{x,T}=\frac{\partial^{2}\zeta}{\partial x\partial T}$. The subscript $p$ indicates that all derivatives are evaluated at the point $(x_{p},T_{p})$. Symbolically, the derivative of $\zeta$ with respect to a generic parameter (such as $x$ or $T$) can be expressed in terms of $\bar{w}$, $\sigma$, and $w_{0}$ as follows
@@ 3.1.1 Pseudo-critical temperature
In the regime where the off-diagonal coupling is small $(0<w_{0}\ll w_{a}\sim w_{b}$), the condition for the pseudo-critical temperature follows from the general minimum conditions for the spectral gap $\zeta(x,T)$, as derived in Appendix 7. In particular, Eq. (74) gives the temperature minimization of $\zeta$, and Eq. (79) provides control parameter minimization. In the zeroth-order limit where excited-state contributions are negligible and $w_{0}\to 0$, these lead directly to the simplified balance condition
@@
A pseudotransition occurs when the dominant eigenvalues $w_{a}$ and $w_{b}$ of two weakly coupled sectors become nearly degenerate. Assuming the off-diagonal coupling $w_{0}$ is small, the pseudo-critical temperature $T_{p}$ can be estimated using Eq.  (20), derived in Appendix  7. Retaining only the ground and first excited states in each sector, one obtains:
@@
where $g_{\alpha,0}$ ($g_{\alpha,1}$) denote the degeneracies of the ground (first excited) states in sector $\alpha\in\{a,b\}$, respectively. The quantity $\epsilon_{p}=\varepsilon_{a,0}(x_{p})-\varepsilon_{b,0}(x_{p})$ is the energy difference between the ground states, and $\delta_{a,1}(x_{p})$, $\delta_{b,1}(x_{p})$ are the excitation gaps within each sector. If excited states are negligible, this expression reduces to the simpler form
@@
which reproduces known results used in various studies of this anomalous behavior [12] This confirms that the sharpness and position of the crossover are governed by both ground-state degeneracies and excitation structures. At $T_{p}$, the statistical weights of the two dominant sectors become comparable, resulting in a sharp but analytic thermodynamic crossover.
3.2 Interface residual entropy
At zero temperature, the entropy of a system reflects the macroscopic degeneracy of its ground state. When a control parameter $x$ is tuned to a critical value $x_{c}$, the system may reach a phase boundary where two distinct ground-state configurations coexist. In such cases, both sectors, labeled $a$ and $b$, contribute significantly to the partition function, giving rise to a finite entropy even at $T=0$. This entropy, associated with the interface between competing ground states, is known as the critical residual entropy.
At finite temperature $T_{p}$ where a pseudotransition occurs, the statistical weights of both sectors become comparable, leading to the condition
@@ 3.2 Interface residual entropy
where $\bar{\varepsilon}=\frac{\varepsilon_{a,0}(x)+\varepsilon_{b,0}(x)}{2}$ and $\epsilon=\varepsilon_{a,0}(x)-\varepsilon_{b,0}(x)$.
Using the condition given by (22) we have $\frac{g_{a,0}}{g_{b,0}}=e^{\epsilon/T_{p}}$, which allows us to express the average weight as
@@
The corresponding residual entropy at the phase boundary is then
$\mathcal{S}_{0}^{\mathrm{c}}=\ln\left(\sqrt{g_{a,0}g_{b,0}}\right).$ (25)
At the critical point $x=x_{c}$, where the ground-state energies of the dominant sectors coincide, $\varepsilon_{a,0}(x_{c})=\varepsilon_{b,0}(x_{c})=\bar{\varepsilon}_{c}$, this result reflects a situation in which both sectors contribute symmetrically to the thermodynamics.
This entropy corresponds to the geometric mean of the ground-state degeneracies and signals a dynamically balanced configuration, enforced by the pseudotransition condition of Eq. (22). It differs from a simple statistical mixture: the near-degeneracy and suppressed coupling lead to coherent coexistence rather than thermodynamic phase separation.
@@ 6 Conclusion
In summary, pseudotransitions originate not from broken analyticity, but from the interplay of spectral proximity and suppressed coupling within irreducible transfer matrices. Nearly block-diagonal irreducibility thus provides a natural explanation for sharp thermodynamic responses in 1D systems, without contradicting the general absence of phase transitions[3].
@@ Acknowledgments
This work was partially supported by the Brazilian funding agencies CNPq and FAPEMIG.
7 Derivation of pseudo-critical temperature
This appendix provides the algebraic derivation of the pseudo-critical conditions, based on the spectral structure of a two-sector transfer matrix and the behavior of the spectral gap $\zeta(x,T)$.
For simplicity, we assume that each effective matrix element is given by a Boltzmann sum over internal excitations. Assuming the system is gapped, we write,
@@ 7 Derivation of pseudo-critical temperature
where $\hat{g}_{\nu,k}=\frac{g_{\nu,k}}{g_{\nu,0}}$ representing the relative degeneracy of the $k$-th excited level, and $\delta_{\nu,k}(x)=\varepsilon_{\nu,k}(x)-\varepsilon_{\nu,0}(x)$ being the excitation gap. In the low-temperature limit, we approximate $\eta_{\nu}(x,T)\approx 1$, neglecting excited-state contributions.
The temperature $T$ and $x$-derivatives of $w_{\nu}$ then follow
@@
Let us define the mean and difference of ground-state energies as $\bar{\varepsilon}=\frac{\varepsilon_{a,0}(x)+\varepsilon_{b,0}(x)}{2}$ and $\epsilon=\varepsilon_{a,0}(x)-\varepsilon_{b,0}(x)$, both of which depend on $x$. Then, using the notation in the main text and the result from (69), we compute

### v2

@@ 1 Introduction
Despite these advances, most analyses describe the sharp crossovers without explaining their origin. Here we identify the generic two-sector spectral structure that produces pseudotransitions, derive the corresponding pseudo-critical condition, and show how an interface residual entropy arises from the constrained low-temperature limit along the pseudo-critical line. This yields a unified perspective on the mechanism behind the anomalous behavior reported in many one-dimensional models.
@@ 2.2 Block-diagonal reducible matrices
A prominent example is the Kittel model [43], in which the local constraints forbid transitions between closed and open configurations, so the transfer matrix remains reducible at all finite temperatures. Similar behavior appears in $q$-state models such as Refs. [25, 45] when the limit of infinite $q$ is taken.
@@
This simple two-block decomposition captures the mechanism by which strict reducibility allows genuine thermodynamic singularities. The Kittel model[43] is a representative example of this situation. We return to this point in Sec. 2.3, where the comparison with nearly reducible but fully connected matrices clarifies why pseudotransitions arise even though the transfer matrix remains irreducible.
@@ 2.3 Nearly block-diagonal irreducible matrices
$\mathbf{V}=\begin{pmatrix}\mathbf{A}&\mathbf{C}\\ \mathbf{D}&\mathbf{B}\end{pmatrix}$, (7)
where $\mathbf{A}$ and $\mathbf{B}$ are dominant irreducible blocks, while $\mathbf{C}$ and $\mathbf{D}$ contain small off-diagonal couplings. Although $\mathbf{V}$ is formally irreducible, the elements of $\mathbf{C}$ and $\mathbf{D}$ may become suppressed at a given parameter value $x_{p}$, nearly decoupling the sectors $\mathbf{A}$ and $\mathbf{B}$ at a fixed temperature.
@@
$\mathcal{V}_{t}=\begin{pmatrix}w_{a}&\langle\phi_{a}|\mathbf{C}|\psi_{b}\rangle\\ \langle\phi_{b}|\mathbf{D}|\psi_{a}\rangle&w_{b}\end{pmatrix}$, (8)
where $\langle\phi_{a}|\mathbf{C}|\psi_{b}\rangle>0$ and $\langle\phi_{b}|\mathbf{D}|\psi_{a}\rangle>0$, according to the condition of matrix non-negative and irreducibility, then let us define $\langle\phi_{b}|\mathbf{D}|\psi_{a}\rangle\langle\phi_{a}|\mathbf{C}|\psi_{b}\rangle=w_{0}^{2}$.
The emergence of an effective two-sector description follows from the low-temperature spectral structure of the transfer matrix. In strictly reducible cases such as the Kittel model [3, 43], the dominant eigenvalues $w_{a}$ and $w_{b}$ belong to disconnected blocks. In the irreducible but nearly block-structured situation considered here, these quantities play the role of diagonal statistical weights for two competing low-energy configurations, while the off-diagonal term $w_{0}$ represents the minimal excitation linking them. Near a pseudotransition the dominant eigenvalues lie exponentially closer to each other than to the rest of the spectrum, so subdominant eigenvectors decay on much shorter length scales and remain almost orthogonal to the space spanned by $|\psi_{a}\rangle$ and $|\psi_{b}\rangle$. Since $w_{0}$ is also exponentially small, hybridization with higher sectors is negligible, and the thermodynamics is effectively governed by the resulting $2\times 2$ structure.
Thus, the eigenvalues of $\mathcal{V}_{t}$ are:
@@ 3.1 Expansion of the spectral gap and pseudo-critical point
To analyze the thermodynamic behavior close to a pseudotransition, we expand the normalized spectral gap $\zeta(x,T)$ in a Taylor series around the anomalous point $(x_{p},T_{p})$. Analyticity of $\zeta$ in both variables is guaranteed for any finite irreducible transfer matrix by Kato’s theorem [42], and the presence of the mixing term $w_{0}>0$ ensures that $\zeta$ remains strictly positive even when $w_{a}=w_{b}$. The detailed expressions for the derivatives, including mixed terms, are derived in Appendix A.
For small deviations $\delta T=T-T_{p}$ and $\delta x=x-x_{p}$, the spectral gap admits the local expansion
@@
where all derivatives ($\zeta_{T}=\frac{\partial\zeta}{\partial T}$, $\zeta_{x}=\frac{\partial\zeta}{\partial x}$, $\zeta_{T^{2}}=\frac{\partial^{2}\zeta}{\partial T^{2}}$, $\zeta_{xx}=\frac{\partial^{2}\zeta}{\partial x^{2}}$, $\zeta_{x,T}=\frac{\partial^{2}\zeta}{\partial x\partial T}$) are evaluated at $(x_{p},T_{p})$. This expansion is valid within the narrow region where the pseudo-critical crossover occurs, that is, for $|\delta T|$ and $|\delta x|$ small enough that higher-order terms remain negligible. The mixed derivative $\zeta_{xT}$ quantifies how variations in the control parameter and temperature jointly affect the curvature of the gap valley, determining the tilt of the minimum in the $(x,T)$ plane and how its position shifts when $x$ and $T$ vary simultaneously.
Symbolically, the derivative of $\zeta$ with respect to a generic parameter (such as $x$ or $T$) can be expressed in terms of $\bar{w}$, $\sigma$, and $w_{0}$ as follows
@@ 3.1.1 Pseudo-critical temperature
In the regime where the off-diagonal coupling is small $(0<w_{0}\ll w_{a}\sim w_{b}$), the condition for the pseudo-critical temperature follows from the general minimum conditions for the spectral gap $\zeta(x,T)$, as derived in Appendix A. In particular, Eq. (74) gives the temperature minimization of $\zeta$, and Eq. (79) provides control parameter minimization. In the zeroth-order limit where excited-state contributions are negligible and $w_{0}\to 0$, these lead directly to the simplified balance condition
@@
Physically, the condition $w_{a}=w_{b}$ implies equality of the free energies of two sectors, kind of true phase transition when ignoring $w_{0}$ contribution. However, when $w_{a}\approx w_{b}$, the two leading eigenvalues (11) is out of the minimum and sharp changes vanishes.
A pseudotransition occurs when the dominant eigenvalues $w_{a}$ and $w_{b}$ of two weakly coupled sectors become nearly degenerate. Assuming the off-diagonal coupling $w_{0}$ is small, the pseudo-critical temperature $T_{p}$ can be estimated using Eq.  (20), derived in Appendix  A. Retaining only the ground and first excited states in each sector, one obtains:
@@
where $g_{\alpha,0}$ ($g_{\alpha,1}$) denote the degeneracies of the ground (first excited) states in sector $\alpha\in\{a,b\}$, respectively. The quantity $\epsilon_{p}=\varepsilon_{a,0}(x_{p})-\varepsilon_{b,0}(x_{p})$ is the energy difference between the ground states, and $\delta_{a,1}(x_{p})$, $\delta_{b,1}(x_{p})$ are the excitation gaps within each sector.
If excited states are negligible, this reduces to
@@
a widely used expression in the literature[12]. A positive $T_{p}$ occurs only when energetic and entropic contributions compete: $\epsilon_{p}<0$ with $g_{a,0}<g_{b,0}$, or $\epsilon_{p}>0$ with $g_{a,0}>g_{b,0}$. Otherwise, one sector dominates for all temperatures and no pseudotransition appears in the canonical ensemble.
The degeneracies determine where the condition $w_{a}=w_{b}$ is met, since the entropic terms $T\ln g_{\alpha,0}$ shift the effective free energies of the two sectors. A sector with larger ground-state multiplicity may therefore dominate even when its energy is higher, as explored in Sec. 4. When the degeneracies are equal this entropic contribution cancels, and the balance condition is determined by the structure of the low-lying excitations and $T_{p}$ is obtained from (21), as discussed in Sec. 5. Equal degeneracies do not guarantee a pseudotransition; such behavior occurs only when the excitation gaps differ sufficiently to produce an exponentially sharp, yet analytic, crossover.
3.2 Zero temperature limit along the pseudo-critical line
We consider the low-temperature limit restricted to the pseudo-critical line $T=T_{p}(x)$, defined in the previous subsection by the condition $w_{a}(x,T_{p})=w_{b}(x,T_{p})$. The diagonal weights consist of an energetic and an entropic factor: $\varepsilon_{\alpha,0}(x)$ is the ground-state energy of sector $\alpha$, while $g_{\alpha,0}$ counts the number of microstates compatible with that energy. Using their low-temperature form, the balance condition becomes
@@ 3.2 Zero temperature limit along the pseudo-critical line
where $\bar{\varepsilon}=\frac{\varepsilon_{a,0}(x)+\varepsilon_{b,0}(x)}{2}$ is the average ground-state energy and $\epsilon=\varepsilon_{a,0}(x)-\varepsilon_{b,0}(x)$ is their difference.
Rearranging yields the constrained relation $\frac{g_{a,0}}{g_{b,0}}=e^{\epsilon/T_{p}}$, which allows us to express the average weight as
@@
The entropy per site associated with this balanced configuration is therefore
$\mathcal{S}_{0}^{\mathrm{c}}=\ln\left(\sqrt{g_{a,0}g_{b,0}}\right),$ (25)
which is the geometric mean of the ground-state degeneracies. An alternative rigorous demonstration of $\mathcal{S}_{0}^{c}$ is given in Appendix B.
The entropy (25) is valid when the low-temperature behavior is controlled by two discrete sectors with finite excitation gaps, so that their asymptotic weights remain accurate along the pseudo-critical line. In this regime the off-diagonal mixing is exponentially small, and the condition $w_{a}(T_{p})=w_{b}(T_{p})$ determines the leading contribution to the dominant eigenvalue. The limit $T\to 0$ must be taken along this line; at fixed parameters one sector always dominates and the entropy reduces to $\ln\left(g_{\alpha,0}\right)$. If additional sectors compete, or if the mixing becomes comparable to the diagonal weights, the two-sector description fails and Eq. (25) no longer applies.
The residual entropy assumes an intermediate value because irreducibility ($w_{0}>0$) enforces strictly positive statistical weights for both sectors along the pseudo-critical line. When the limit $T\to 0$ is taken with $w_{0}>0$, the stationary measure retains contributions from both sectors and yields $\mathcal{S}_{0}^{c}=\ln\sqrt{g_{a,0}g_{b,0}}$. In contrast, imposing $w_{0}=0$ before the low-temperature limit makes the transfer matrix reducible, causing the system to collapse onto a single dominant sector and producing $\max(\ln g_{a,0},\ln g_{b,0})$. The two limits do not commute, and the maximal value arises only in the strictly reducible case.
@@ 5.0.2 Core-ferromagnetic (CFM) phase
Before deriving the effective thermodynamics, we summarize the structure of the transfer matrix. The full $1458\times 1458$ matrix decomposes into invariant subspaces fixed by the $C_{6v}$ symmetry of the hexagonal shell and by the allowed configurations of the core spin, as detailed in Sec. 5.1. This yields a finite set of symmetry-related blocks, the largest being $92\times 92$ as detailed in Ref. [38]. Near the pseudo-critical region, a single block dominates and becomes effectively two-dimensional due to the competition between two low-energy sectors as detailed in Sec.5.2. The resulting $2\times 2$ matrix reproduces the leading eigenvalues shown in Fig. 5, confirming the reduction.
@@ 6 Conclusion
In summary, pseudotransitions originate not from broken analyticity, but from the interplay of spectral proximity and suppressed coupling within irreducible transfer matrices. Nearly block-diagonal irreducibility thus provides a natural explanation for sharp thermodynamic responses in 1D systems, without contradicting the general absence of phase transitions[3]. The present analysis is restricted to discrete energy spectra, and the extension to continuous spectra will be addressed elsewhere.
@@ Acknowledgments
This work was partially supported by the Brazilian agencies CNPq and FAPEMIG.
Appendix A Derivation of pseudo-critical temperature
This appendix presents the algebraic derivation of the pseudo-critical conditions using the discrete spectral structure of a two-sector transfer matrix and the properties of the spectral gap $\zeta(x,T)$. These assumptions ensure that the low-temperature asymptotics of the sector weights remain valid.
For each sector we write the effective matrix element as a Boltzmann sum over internal excitations,
@@ Appendix A Derivation of pseudo-critical temperature
collects the contributions of excited states. Here $\hat{g}_{\nu,k}=\frac{g_{\nu,k}}{g_{\nu,0}}$ denotes the relative degeneracy of the $k$-th excited level, and $\delta_{\nu,k}(x)=\varepsilon_{\nu,k}(x)-\varepsilon_{\nu,0}(x)$ being the excitation gap.
In the low-temperature regime $T\ll\delta_{\nu,1}$, all excited states are exponentially suppressed and the correction factor satisfies $\eta_{\nu}(x,T)\approx 1$. Under these conditions the dominant sector weights reduce to their ground-state contributions, and the pseudo-critical condition can be derived in closed form.
The temperature $T$ and $x$-derivatives of $w_{\nu}$ become
@@
Here we assume the mean and difference of ground-state energies as $\bar{\varepsilon}=\frac{\varepsilon_{a,0}(x)+\varepsilon_{b,0}(x)}{2}$ and $\epsilon=\varepsilon_{a,0}(x)-\varepsilon_{b,0}(x)$, both of which depend on $x$. Using the notation in the main text and the result from (69), we compute
@@
In summary, Eq. (67) and the resulting formulas hold in the regime $T\ll\delta_{a,1},\delta_{b,1}$ and $w_{0}\ll w_{a},w_{b}$, that is, within the two-sector low-temperature window where pseudotransitions arise. The approach assumes a discrete local spectrum and does not extend to continuous or gapless cases, for which the hierarchy of Boltzmann weights assumed here no longer holds.
Appendix B Entropy at the pseudo-critical point
This appendix provides a compact derivation of the entropy evaluated at the pseudo-critical temperature $T_{p}$. All quantities and expressions introduced in Sec. 3 (sector weights, derivatives, eigenvalues $\lambda_{1,2}$, and the low-temperature forms) are used here.
B.1 Derivative of the dominant eigenvalue
Differentiating the largest $\lambda_{1}(T)$ obtained from (9) and then imposing the balance condition $w_{a}(T_{p})=w_{b}(T_{p})$ yields
$\lambda^{\prime}_{1}(T_{p})=\frac{w^{\prime}_{a}(T_{p})+w^{\prime}_{b}(T_{p})}{2}+w^{\prime}_{0}(T_{p}).$ (80)
Along the pseudo-critical line, the diagonal weights satisfy the Eq.(69) we have $w_{\nu}^{\prime}(T)=w_{\nu}(T)\frac{\varepsilon_{\nu,0}}{T^{2}},$so that
$\frac{w^{\prime}_{a}(T_{p})+w^{\prime}_{b}(T_{p})}{2}=\frac{w(T_{p})}{2T_{p}^{2}}\bigl(\varepsilon_{a,0}+\varepsilon_{b,0}\bigr),$ (81)
where $w(T_{p})=w_{a}(T_{p})=w_{b}(T_{p})$. The off-diagonal weight has the low-temperature form $w_{0}(T)\sim\exp[-\varepsilon_{0,0}/T]$, hence using (69) we have $w^{\prime}_{0}(T_{p})=w_{0}(T_{p})\frac{\varepsilon_{0,0}}{T_{p}^{2}}$. The ratio of the off-diagonal derivative to the diagonal one is therefore
$\frac{w^{\prime}_{0}(T_{p})}{w^{\prime}_{a}(T_{p})+w^{\prime}_{b}(T_{p})}=\frac{w_{0}(T_{p})}{w(T_{p})}\cdot\frac{\varepsilon_{0,0}}{\varepsilon_{a,0}+\varepsilon_{b,0}}.$ (82)
Because pseudo-transition behavior requires $w_{0}(T_{p})\ll w(T_{p})$, and the energy ratio is $\mathcal{O}(1)$, so the expression in (82) is exponentially small. Hence
$w^{\prime}_{0}(T_{p})\ll w^{\prime}_{a}(T_{p})+w^{\prime}_{b}(T_{p}),$ (83)
and to leading order
$\lambda^{\prime}_{1}(T_{p})\simeq\frac{w(T_{p})}{2T_{p}^{2}}\bigl(\varepsilon_{a,0}+\varepsilon_{b,0}\bigr).$ (84)
B.2 Entropy at the pseudo-critical point
Using
$\mathcal{S}(T)=\ln\lambda_{1}(T)+T\frac{\lambda^{\prime}_{1}(T)}{\lambda_{1}(T)},$ (85)
and $\lambda_{1}(T_{p})=w(T_{p})+\mathcal{O}(w_{0})$, we obtain
$\mathcal{S}(T_{p})\simeq\ln w(T_{p})+\frac{\varepsilon_{a,0}+\varepsilon_{b,0}}{2T_{p}}.$ (86)
At low temperature the dominant weight satisfies
$\ln w(T_{p})=\ln g_{a,0}-\frac{\varepsilon_{a,0}}{T_{p}},$ (87)
substituting into (86) gives
$\mathcal{S}(T_{p})=\ln g_{a,0}+\frac{\varepsilon_{b,0}-\varepsilon_{a,0}}{2T_{p}}.$ (88)
The balance condition at $T_{p}$ yields
$\frac{\varepsilon_{b,0}-\varepsilon_{a,0}}{T_{p}}=\ln g_{b,0}-\ln g_{a,0},$ (89)
and substituting (89) into (88) leads directly to
$\mathcal{S}(T_{p})=\frac{1}{2}\bigl(\ln g_{a,0}+\ln g_{b,0}\bigr).$ (90)
This equals the interface residual entropy,
$\mathcal{S}_{0}^{c}=\ln\sqrt{g_{a,0}g_{b,0}}.$ (91)
These results hold when the low-temperature behavior is governed by two discrete sectors with $w_{0}\ll w_{a},w_{b}$ and when the limit is taken along the pseudo-critical line $T_{p}(x)$; if additional sectors become competitive or the spectrum is not gapped, the derivation does not apply.